OpenStudy (malaya):

Which of the following is an equation of the line that passes through (3,-5) with a slope of -3? A . y - 5 = -3(x - 3) B . y + 5 = -3(x - 3) C . y - 5 = 3(x + 3) D . y + 5 = -3(x + 3)

OpenStudy (ryan123345):

x1=3, y1=-5, m=-3. The formula is y-y1=m(x-x1). When you plug the numbers in youll get y-(-5)=-3(x-3). Two negatives give you a positive so you get y+5=-3(x-3)